OpenStudy (anonymous):

Simplify the following expression. 4x − 3y + 2z − 3x − y + 3z = ?

OpenStudy (anonymous):

4x - 3y + 2z - 3x - y + 3z = x - 4y + 5z.

OpenStudy (anonymous):

group the like terms. (4x-3x)+(-3y-y)+(2z+3z) =x-4y+5z <-- simplify

OpenStudy (anonymous):

4x − 3y + 2z − 3x − y + 3z =(4-3)x+(-3-1)y+(2+3)z =x-4y+5z

OpenStudy (anonymous):

The solution to this is simple, combine like terms like this: 4x-3x= x -3y-y= -4y 2z+3z= 5z There for the answer is x-4y+5z